Write -1/2 + 2/3 - 3/4 + 4/5 -... to n terms in sigma notation.
1
Expert's answer
2020-04-30T19:59:18-0400

"-\frac{1}{2}+\frac{2}{3}-\frac{3}{4}+\frac{4}{5}-...= \sum_1^n(-1)^r\frac{r}{r+1}"